DTL RF conditioning
RF conditioning of Drift Tube Linacs 2-4 is in progress in the normal conducting linac. RF power is supplied by the klystrons in the normal conducting gallery at slowly increasing capacity. The goal is to meet the final nominal power level of 14 Hz, 3.2 ms.
The operations are supervised from the ESS main control room.
The RF conditioning is a crucial step in making the normal conducting linac ready for Beam Commissioning later this spring.
RF power sources
With the modulator for cell 190B moved to its permanent place, all nine modulators powering the medium beta gallery are now in position in RF cells 150-190. All RF racks in the medium beta section has been energised, and power tests are ongoing in both the spoke and medium beta gallery.
Rack installations are ongoing in the high beta section.
Cryomodule Test Stand
A cleanroom is in place to perform repairs of medium beta cryomodule CM02, on which a vacuum leak was detected across the ceramic window of a coupler. The damaged part will be carefully exchanged, under supervision by the French in-kind supplier, CEA.
Drift Tube Linac 5, the last part of the DTL section which is to be installed in the tunnel after the upcoming beam commissioning up to DTL4 has been completed, is under preparation in the DTL workshop. Module assembly is under way, and drift tube installations are performed inside a cleanroom confinement.
Monolith inner shielding
This week, the Reflector base shielding block was successfully re-installed into the Monolith vessel. After defects were detected both in regards to cleanliness and the quality of the welds, several months of defining a repair strategy and implementation of corrective measures followed. This work resulted in an approved multi-step method for flushing the interior of the piping system and an advanced re-welding process.
With the first block in place and all procedures running smoothly, a ramp up in inner shielding installation is expected over the next months.
Target Test Stand
Preparations to install the moderator/reflector plug next to the Target wheel assembly inside the Target mock-up and test stand is now under way. The equipment will be lowered into the bucket adjacent to the wheel. Functional synchronised testing will then be performed inside the test stand.
Neutron Beam Port Inserts (NBPI)
Optics have been installed inside the Neutron Beam Port Insert for C-SPEC. Vacuum tests will follow.
With the Target Monolith inner shielding installation now back on track, preparations to start the north sector installation of Neutron Beam Port Inserts and Plugs (NBPI and NBPP) as well as Neutron Beam Windows, are under way.
Installation of the section 2 BEER guide, which comprises 64.5 metres of neutron guide integrated into vacuum housing, is completed. Final leak tests and final acceptance tests are now to be held.
Section 1 BEER guide, starting a few metres from the west bunker wall, will follow in a couple of weeks. This section is 48.6 meters long.